Cheetahs are considered as the fastest land animal. One cheetah observed in the

wild started running from rest and increased its velocity uniformly until it reached

99 km/h. It covered 40 m while running. What is its acceleration? How much time

will it take to cover the said displacement?


1
Expert's answer
2021-04-12T06:56:32-0400

By definition, the acceleration is:


"a = \dfrac{v_f - v_i}{t} = \dfrac{v_f}{t}"

where "v_f = 99km\/h = 27.5 m\/s" is the final speed, "v_i = 0" is the initial speed, and "t" is the time of accelration.

The displacement is given as follows:


"d =v_it + \dfrac{a t^2}{2} = \dfrac{a t^2}{2}"

Substituting the expression for "a" and "d = 40m" and expressing "t", obtain:


"d = \dfrac{v_ft}{2}\\
t = \dfrac{2d}{v_f} = \dfrac{2\cdot 40m}{27.5m/s} \approx 2.9s"

Substituting the expression for "t" into the formula for "a", obtain:


"a = \dfrac{v_f}{\dfrac{2d}{v_f}} = \dfrac{v_f^2}{2d}\\
a = \dfrac{(27.5m/s)^2}{2\cdot 40m} \approx 9.5m/s^2"

Answer. The acceleration is 9.5 m/s^2, the time is 2.9s.